OpenStudy (anonymous):

B is between A and C. Find BC if AB = 3, BC = 4x + 1, AC = 8. A.1 B.5 C.-5 D.-1

OpenStudy (seamus58):

AB=3,BC=4x+1andAC=8. 3+4x+1=8 4+4x=8 4x=4 x=1 BC=4x+1 if x=1, then BC 4*1+1=5 The answer is B.5